# Verifying Device Drivers with Pancake

### v1

@@ 6.2 Future Verification Efforts
6.2.1 Transpiler
Work towards a formal correctness proof for the implementation of the transpiler described in Section 3 is still ongoing. Though the Viper IL code is obtained by the transpiler in a straightforward manner from the original Pancake code, presently we have no formal proof of a semantic correspondence between the input Pancake code and the resulting Viper IL code. Consequently, the transpiler is currently included as part of the trusted computing base, as discussed in Section 6.1.
@@ 6.2.1 Transpiler
Such a proof would complete our end-to-end verification story, by allowing us to soundly infer the correctness of our initial Pancake code from a successful verification run of the corresponding Viper IL code. This assumes the correct operation of the Viper toolchain which our verification step relies on. Existing efforts to validate parts of this toolchain can be found in [Parthasarathy et al., 2024; Gössi, 2016].
@@
Our eventual plan is to complete a mechanised correctness proof for the Pancake to Viper IL transpiler within the HOL4 theorem prover. As a first step towards this, we require formal descriptions of the transpiler, and the semantics of Pancake and the Viper IL. Though the current Rust implementation of the transpiler contains 7 kLOC, a significant portion of this code consists of workarounds which extract the underlying abstract syntax from an unparsed Pancake program without interfacing directly with the existing Pancake parser. However, as part of the CakeML project, the Pancake language already has its parser defined in HOL4. Additionally, the formal syntax and semantics of Pancake have been mechanised in HOL4. Thus, an implementation in HOL4 can access the abstract syntax directly, which could simplify the implementation considerably.
6.2.2 Interactive Theorem Proving for Pancake programs
The Pancake language semantics are defined in the HOL4 interactive theorem prover (ITP). It would be natural to verify a Pancake program directly in HOL4, proving that its semantics agree with its specification: it would significantly shrink the TCB. Furthermore, it would be possible to formally compose the source-level verification with the compiler correctness proof for Pancake (also in HOL4) to give an end-to-end proof that the compiled binary is correct.
Hoare logic
We have defined a Hoare logic for Pancake programs in HOL4. This is a standard approach to verifying programs in an interactive language, and resembles the proof approach of Viper. Programs are decorated with preconditions, postconditions, and loop invariants. The logic specifies how a proof of a precondition/postcondition pair for a function body can be decomposed into proofs about each statement. We have defined the logic in HOL4 and proved its soundness against the Pancake semantics. We have also added an automated proof tactic that drives the logic, doing all the decomposition and re-composition steps, leaving the user to prove verification conditions such as that the precondition they supplied is sufficient, that their loop invariant implies itself at the next iteration, etc.
At the time of writing, this Hoare logic automation makes the proof process far more productive than a naive manual approach. But our proof productivity with this automation is still far short of what is achieved with our Viper-based approach. We think that there is still substantial room for improvement in future work.
One interesting possibility is to try to maximise compatibility between the HOL4 Hoare logic and the fragment of Viper’s annotation syntax we are using in this work. The aim would be that an initial verification in Viper could later be minimally converted into a HOL4 verification, and replayed without the involvement of Viper or the transpiler. This is a potential alternative to verifying the transpiler and would also take Viper and the SMT solvers out of the TCB.
Interaction trees semantics
As mentioned in Section 2.2, Pancake programs that communicate with the outside world are parameterised on a function that models how the outside world behaves. These models are deterministic, which simplifies definitions and compiler proofs but makes modelling realistic devices awkward: from a programmer’s point of view, devices are non-deterministic. Moreover, requiring device models up-front makes it difficult to decouple reasoning about code and about devices.
A more promising approach is interaction trees [Xia et al., 2020], which represent the behaviour of a program as a coinductive tree, with nodes for actions and branches for possible environmental responses. This model naturally accommodates non-determinism, in contrast to the linear-time view of a functional semantics.
We have developed an interaction tree semantics for Pancake and are currently proving its correspondence to the functional semantics. We are also currently applying it to verification of device drivers similar to those described here.
@@ 7 Related Work
A number of prior works have investigated verification of device drivers at various programming language levels. Within this space, researchers have employed different verification approaches like model checking [Kim et al., 2008] and interactive theorem proving [Möre, 2021]. Among these efforts, [Penninckx et al., 2012] developed their verification using VeriFast [Jacobs et al., 2010], a deductive verification approach similar to ours. They also notably extended their analysis to include concurrency properties beyond our current scope. However, a common limitation across all these approaches was the significant gap between the analysed model and the actual executable code—a gap which we narrow by using Pancake’s verified compilation stack.
@@
Finally, there has been also been work on driver synthesis by [Ryzhyk et al., 2009b, 2014] that took as input detailed specifications of interfaces for (1) the device class the driver needs to implement, the (2) device itself and (3) OS service it needs to provide to the rest of the OS, written in a custom specification language. Although, like in our work, their device interface included details such as valid registers and their sizes, it also included more detailed elements like a state transition diagram. In our work, we have a model of device state that we use for specifying and verifying the maintenance of invariants (the valid_device(device) assertion seen in 3 and explained in Section 4.1))—this could in future form the basis for more detailed, state machine-based specifications of internal device states. Note, however, that this synthesis work could not deal with DMA.
@@ Acknowledgements
The development of Pancake was made possible through the generous support of multiple organisations: the UAE Technology Innovation Institute (TII) under the Secure, High-Performance Device Virtualisation for seL4 project, DARPA under prime contract FA 8750-24-9-1000, and UK’s National Cyber Security Centre (NCSC) under project NSC-1686. Finally, Junming Zhao gratefully acknowledges the topup scholarship donated by the winners of the 2023 ACM Software System Award.

### v2

@@ 6.2 Future Verification Efforts
We are continuing to implement more drivers for LionsOS in Pancake and to apply the verification method presented in this paper to those drivers firstly to produce more verified drivers but also to assess and improve the verification process. We also have various verification related future projects based on Pancake, including those we discuss below.
Transpiler verification
@@
Proving the formal correctness of the transpiler implementation, as mentioned in Section 3, is in progress. This means that the transpiler is currently included as part of the TCB, as discussed in Section 6.1. What we aim to prove here is to establish the validity at the level of the Pancake input of the statements proved on at the level of the transpiled Viper IL code. Such a proof would complete our end-to-end verification story, by allowing us to soundly infer the correctness of our initial Pancake code from a successful verification run of the corresponding Viper IL code. This still assumes the correct operation of the Viper toolchain which our verification step relies on. Efforts exist to validate parts of this toolchain [Parthasarathy et al., 2024; Gössi, 2016].
@@
We will do this proof using the theorem prover HOL4, which Pancake compiler is implemented. Though the current Rust implementation of the transpiler contains 7 kLOC, a significant portion of this code consists of workarounds for extracting the underlying abstract syntax from an unparsed Pancake program without interfacing directly with the existing Pancake parser. We can remove these workarounds entirely by connecting the transpilation to the Pancake parser and handling its formal syntax and the semantics directly, which will simplify the implementation considerably, making the verification easier.
Interactive theorem proving for Pancake programs
The formal semantics for Pancake and the compiler for it are developed in the HOL4 interactive theorem prover. It would be natural to verify a Pancake program entirely in HOL4. We have performed a number of small experiments in this style, verifying simple example programs such as a loop that sums the elements of an array. These experiments are an important validation of the Pancake formal semantics. The compiler correctness proof establishes that the formal semantics captures the behaviour of the binary, but that is only useful if it is possible to reason at the level of the semantics.
We have done some of these proofs by directly manipulating the Pancake semantics in HOL4. We have also constructed a Hoare-style precondition-postcondition logic for Pancake programs, which is a standard approach to verification of programs in an imperative language such as Pancake. This gives us a comparison point to our Viper-based approach. Our Hoare logic is far more user-friendly than directly working with the semantics, but far less user-friendly than using the transpiler. One advantage of the Hoare logic is that its correctness results compose directly with the compiler correctness theorem in HOL4, producing a binary correctness theorem with a minimal TCB, useful for small safety-critical programs whose correctness must be certified.
Another strength of HOL4 is its general logic, which allows us to phrase and prove properties that we could not easily express in Viper. We are interested in reasoning about concurrent scenarios. LionsOS encourages systems to be composed from smaller components (such as our Ethernet driver), most of which are internally single-threaded but execute in parallel. We think that studying a parallel protocol in HOL4 and proving correctness of each Pancake component against that protocol using Viper is an interesting direction for future work.
This composition of components will be correct if each component performs the correct I/O actions. The I/O behaviour of a program can be captured in a single value using interaction trees [Xia et al., 2020]. These are a new approach to program semantics that use an infinite coinductive tree of input and output event nodes to model a program. We have developed an alternative semantics for Pancake using interaction trees, and have experimented with proofs that programs obey an I/O protocol using this semantics. This is a promising approach to verifying, for instance, a whole network sub-system.
@@ 7 Related Work
Among various verification efforts in the OS context, a number of prior works have investigated verification of device drivers at various programming language levels. Within this space, researchers have employed different verification approaches like model checking [Kim et al., 2008] and interactive theorem proving [Möre, 2021]. Among these efforts, [Penninckx et al., 2012] developed their verification using VeriFast [Jacobs et al., 2010], a deductive verification approach similar to ours. They also notably extended their analysis to include concurrency properties beyond our current scope. However, a common limitation across all these approaches was the significant gap between the analysed model and the actual executable code—a gap which we narrow by using Pancake’s verified compilation stack.
@@
The Ironclad verification [Hawblitzel et al., 2014] includes a network driver. They verify assembly code by translating it into Dafny code along with assertions. They argue that verification at the level of assembly avoids issues with the compiler; since we have a verified compiler we prefer to verify the sources that the programmer writes. Like in our work, the Ironclad authors annotate each driver function. They mostly establish memory safety and address a public/private data distinction. There does not appear to be any clear statement of the function of the driver or of functional correctness. The Ironclad paper gives some performance data, showing overheads of up to two orders of magnitude for single requests to their verified versus their unverified system, but no measurements of the driver alone or of the system under load, which would be needed to make a direct comparison to our work. They also note that performance is not their priority, whereas our goal here is to achieve both of high guarantees and performance.
Erbsen et al. [2021] offers an end-to-end story of a simple embedded system that includes an Ethernet driver. However, this system is simple enough that it does not require an operating system in the usual sense. They also report ten-fold overhead as the performance results for the whole system. Again, their work does not provide analysis on the driver by itself, making a direct performance comparison difficult.
@@
Finally, there has been also been work on driver synthesis by [Ryzhyk et al., 2009b, 2014] that took as input detailed specifications of interfaces for (1) the device class the driver needs to implement, the (2) device itself and (3) OS service it needs to provide to the rest of the OS, written in a custom specification language. Although, like in our work, their device interface included details such as valid registers and their sizes, it also included more detailed elements like a state transition diagram. In our work, we have a model of device state that we use for specifying and verifying the maintenance of invariants (the valid_device(device) assertion seen in 3 and explained in Section 4.2)—this could in future form the basis for more detailed, state machine-based specifications of internal device states. Note, however, that this synthesis work could not deal with DMA.
@@ Acknowledgements
The development of Pancake was made possible through the generous support of multiple organisations: the UAE Technology Innovation Institute (TII) under the Secure, High-Performance Device Virtualisation for seL4 project, DARPA under prime contract FA 8750-24-9-1000, UK’s National Cyber Security Centre (NCSC) under project NSC-1686, and the German Agentur für Innovation in der Cybersicherheit GmbH (Cyberagentur). Finally, Junming Zhao gratefully acknowledges the topup scholarship donated by the winners of the 2023 ACM Software System Award.